OpenStudy (anonymous):

can you describe multiplying and dividing signed numbers in the real word? ( in your own words please ) :)

OpenStudy (mathstudent55):

I have $10 in each pocket, and I have 4 pockets. How much money do I have? 4 * $10 = $40 I have $10, and I owe $5. I have +$10. The money I owe can be thought of a negative number. My debt is -$5. The net amount of money I have is +$10 + (-$5) = $10 - $5 = $5 I owe 4 of my friends $5 each. How much money do I have? 4 * (-$5) = -$20 I owe a total of $10 to 5 friends. I owe the same to each one. How much do I owe each friend? -$10/5 = -$2